can someone tell me how to slove this problem please?
4x/x^2+3 - 1=0
Note: x^2+3 is all under 4x...the fraction is then subtracted by 1.
please do help thanks!

Answer by stanbon(75887)

4x/(x^2+3) - 1=0
Rewrite as follows:
4x/(x^2+3)=1
Multiply thry by x^2+3 to get:
4x = x^2+3
Rearrange:
x^2-4x+3=0
(x-3)(x-1)=0
x=3 or x=1
Cheers,
Stan H.